OpenStudy (anonymous):

A point M on a segment with endpoints X(1, -2) and Y(10, 3) partitions the segment in a 4:1 ratio. Find M

OpenStudy (campbell_st):

I start this way so M be the point that divides XY .... I'll assume internal division so then \[\frac{XM}{MY} = \frac{4}{1}\] cross multiply \[XM = 4MY\] so then the formula is so then using the points X and Y 1(1, -2) and 4(10,3) this way you get the right ratios with the right points then its simply a case of \[x = \frac{1 \times 1 + 4 \times 10}{4 + 1}`~~~~y = \frac{1 \times -2 + 4 \times 3}{4 + 1}\] to find M(x, y) hope it helps... if the division is external, make 1 of the ratio values negative.

OpenStudy (anonymous):

alright thanks alot